# Two-Timescale Optimization Framework for Decentralized Linear-Quadratic Optimal Control

### v1

@@ 6 Directly Optimizing $\ell_{0}$-penalty
$\min\limits_{\widetilde{W}}\leavevmode\nobreak$ $u_{1,\sigma}(\widetilde{W},X_{k};\lambda)$ (53)
@@
where $\mathcal{X}_{1}=\{\widetilde{W}\in\mathbb{R}^{p^{2}}\colon\widetilde{W}\in\Gamma_{+}^{p},\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},\widehat{A}\widetilde{W}=0\}$ which is closed convex set. First, we claim that $\{H_{\sigma}(\widetilde{W}_{k})\}$ is monotonically nonincreasing.
@@
Given an initial feasible point $\widetilde{W}_{0}$, let $\widetilde{W}_{k}$ be the iterative sequence generated by Algorithm 4 with $u_{1},u_{2}$ given by (52), (49). Then
$H_{\sigma}(\widetilde{W}_{k+1})\leq H_{\sigma}(\widetilde{W}_{k}).$
Proof.
Let $F(\widetilde{W},X_{k})=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})$. According to the fact that
@@
$H_{\sigma}(\widetilde{W}_{k+1})\leq u_{1,\sigma}(\widetilde{W}_{k+1},X_{k};\lambda)\leq u_{1,\sigma}(\widetilde{W}_{k},X_{k};\lambda)=F(\widetilde{W}_{k},\widetilde{W}_{k})=H_{\sigma}(\widetilde{W}_{k}).$
The proof is complete. ∎
Obviously, $\{H_{\sigma}(\widetilde{W}_{k})\}$ is bounded below by 00. Hence, by basic mathematic analysis, $H_{\sigma}(\widetilde{W}_{k})\downarrow H_{\sigma}^{\infty}$, where $H_{\sigma}^{\infty}$ is a positive constant. An inquiry that arises is whether or not $H_{\sigma}^{\infty}=H_{\sigma}^{*}\doteq\min H_{\sigma}(\widetilde{W})$. Regretfully, the above inquiry remains open, but we have the following result.
@@
Every cluster point $z=(z_{1},z_{2})$ of the iterates generated by the Algorithm 4 with $u_{1},u_{2}$ given by (52), (49) is a coordinatewise minimum of the optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$). In addition, if $H_{\sigma}(\cdot,\cdot)$ is regular at $z$, then $z$ is a stationary point of ($\textbf{P}^{\prime\prime}_{\sigma}$).
Proof.
For a feasible $X_{0}=(\widetilde{W}_{0},y_{0})$, by (47), the sublevel set
@@
is compact. By lemma 5, optimization problem ($\textbf{BP}\sigma-2$) has a unique solution for any point $X_{r-1}\in\mathcal{X}$. Then, the proof is complete by Theorem 2 of [40]. ∎
There exists a significant disparity between coordinatewise minimizer and stationary point (local minimizer). When the objective function is regular at coordinatewise minimizer, the coordinatewise minimum becomes a local minimum by the above theorem. In Lemma 3.1 of [45], there is a few discussion about regularity property of a Gateaux-differentiable function, but $H_{\sigma}(\widetilde{W},y)$ is generally not Gateaux-differentiable in this paper. Furthermore, if we assume that $H_{\sigma}^{\infty}=H_{\sigma}^{*}$ holds, we can obtain the following theorem.
@@
Proof.
Since $f_{\sigma}(|(V_{2}\otimes V_{1})\widetilde{W}|)\geq 0$, it is obvious that, for all $\alpha\in\mathbb{R}$, ${\rm lev}_{\leq\alpha}H_{\sigma}(\widetilde{W})$ is compact. Then the set
@@
$\min_{\widetilde{W}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (54)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
@@
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|\widetilde{P}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (55)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},\leavevmode\nobreak\ i=1,\dots,M,$
@@
$\min_{\widetilde{W},\widetilde{P}}\leavevmode\nobreak\ \leavevmode\nobreak$ $f_{1}(\widetilde{W};\lambda)+f_{2}(\widetilde{W})+g(\widetilde{P})$ (56)
${\rm s.t.}\leavevmode\nobreak\ \leavevmode\nobreak$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
Interestingly, $g(\widetilde{P})$ is nothing but a weighted-$\ell_{1}$ norm with weight values given by $y_{k}$. Hence, compared with optimization problem (11), the only difference is the smooth and strongly convex term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ in $f_{1}(\widetilde{W};\lambda)$. Thus, we can utilize optimization scheme (13) by slightly modifying Algorithm 2 (just change $w_{ij}$ to $y_{k}^{(ij)}$, where $y_{k}^{(ij)}$ is the $(i,j)$-th element of $y_{k}$). The following theorem gives the convergence rate of optimization scheme (13) to solve (56).
Theorem 10.
Under the initial setting
$\gamma_{0}=\frac{1}{\lambda}>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\rm and }\beta_{0}>0$
and the condition
$\left(L_{f}\beta_{k}\theta_{k}+\gamma_{k}\|B\|^{2}\right)\alpha_{k}^{2}=\gamma_{k}\beta_{k}\theta_{k},$
we have $\{\widetilde{W}_{k},\widetilde{P}_{k}\}\subset\mathbb{R}^{p^{2}}\times\mathbb{R}^{mn}$ and
$\left\|\mathcal{A}\widetilde{W}_{k}+\mathcal{B}\widetilde{P}_{k}\right\|\leq\theta_{k}\mathcal{R}_{0},\quad\left|F(\widetilde{W}_{k},\widetilde{P}_{k})-F^{*}\right|\leq\theta_{k}\left(\mathcal{E}_{0}+\left\|\lambda^{*}\right\|\mathcal{R}_{0}\right),$
$\theta_{k}\lesssim\frac{\|\mathcal{B}\|}{\sqrt{\beta_{0}}k}+\min\left\{\frac{1}{k^{2}},\exp\left(-\frac{k}{4}\right)\right\},$
provided that $\gamma_{0}\beta_{0}\leq L_{f}\beta_{0}+\gamma_{0}\left\|\mathcal{B}\right\|^{2}$.
Proof.
By noticing that $f_{1}(\widetilde{W})$ is $\frac{1}{\lambda}$-strongly convex and $\frac{1}{\lambda}$-smooth, this theorem is a direct corollary of Theorem 4.1 of [31]. ∎
@@
1:Given : Feasible initial point $X_{0}=(\widetilde{W}_{0},y_{0})\in\mathcal{X}_{1}\times\mathbb{R}_{+}^{mn},\lambda>0,\gamma>0,\sigma_{0}\leq 1,\alpha\in(0,1)$
2:Result: $X_{k}=(\widetilde{W}_{k},y_{k})$
3:for $i=1,2,3\dots$ do
4:     Set $\sigma_{i}=\alpha\sigma_{i-1}$
5:     Set $k=0$
6:     for $k=0,1,2,...$ do
7:         Let $k=k+1,i=(k\mod 2)+1$
8:         if $i==1$ then
9:              $\widetilde{W}_{k}\leftarrow$ALM based Decentralized LQ
10:              Set $y_{k}=y_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
11:         end if
12:         if $i==0$ then
13:              Update $y_{k}=\nabla f_{\sigma_{i}}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right)$
14:              Set $\widetilde{W}_{k}=\widetilde{W}_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
15:         end if
16:         if Stopping Criterion==absent$=$ =True then
17:              Break
18:         end if
19:     end for
20:     Set $X_{0}=X_{k}$
21:end for
@@
the iterative sequence of the algorithm BSUM can also converge to the coordinatewise minimum of $H_{\sigma}(\widetilde{W},y)$. However, we introduce the proximal term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ into $u_{1,\sigma}$, which appears to be gilding the lily or potentially resulting in a decelerated convergence of the BSUM algorithm. Empirically, the proximal term makes the objective of (56) be strongly convex and accelerate the solving for subproblem (56). Therefore, a trade-off exists between the convergence rate of the algorithm BSUM for optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) and the convergence rate of (13) for subproblem (56).
@@
The algorithm exhibits a significant level of complexity, and there exists no empirical evidence to support the superiority of Algorithm 5 over Algorithm 2 (or Algorithm 3). Therefore, this section mostly focuses on theoretical discourse, particularly about the variational characteristic of the optimization problem induced by the $\ell_{0}$-penalty and the relation between $\ell_{1}$-penalty problem (8) and $\ell_{0}$-penalty problem. It is recommended that, in the pursuit of a decentralized controller, Algorithm 2 (or Algorithm 3) is employed instead of Algorithm 5.
@@ 7 Numerical Examples
$A=\begin{bmatrix}0.2220&0.9186&0.7659\\ 0.8707&0.4884&0.5184\\ 0.2067&0.6117&0.2968\end{bmatrix},B_{1}=I_{3},B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
By Algorithm 2, we can obtain the relation between $\gamma$ and $N$ (the number of 00 in feedback gain $K$; see Figure 2).
@@
Specifically, when $\gamma=10$, we can obtain the following result (see Figure 3).
@@
The optimal $W$ is given by
$W=\begin{bmatrix}2.1348&0&0&1.3084&0\\ 0&0.2698&0&0.6783&0\\ 0&0&1.0833&0&1.4155\\ 1.3084&0.6783&0&2.5604&0.0129\\ 0&0&1.4155&0.0129&1.8711\end{bmatrix},$
and the optimal decentralized feedback gain $K$ is given by
@@
The responses of all the state variables are illustrated in Figure 4, and it can be seen that the stability of system is guaranteed.
@@
$A=\begin{bmatrix}0.3079&0.1879&0.1797&0.2935&0.6537\\ 0.5194&0.2695&0.5388&0.9624&0.5366\\ 0.7683&0.4962&0.2828&0.9132&0.9957\\ 0.7892&0.7391&0.7609&0.5682&0.1420\\ 0.8706&0.1950&0.2697&0.4855&0.9753\end{bmatrix},B_{2}=\begin{bmatrix}0.6196&0.6414\\ 0.7205&0.9233\\ 0.2951&0.8887\\ 0.6001&0.6447\\ 0.7506&0.2956\end{bmatrix}$
@@
Letting $\gamma=10$, and by Algorithm 2, the optimal $W$ is given by
$W=\begin{bmatrix}2.5248&0&0&0&0&3.3422&2.2528\\ 0&3.0027&0&0&0&0&5.6741\\ 0&0&4.2745&0&0&0&4.4252\\ 0&0&0&0.9312&0&0&3.5367\\ 0&0&0&0&0.3125&1.7188&0\\ 3.3422&0&0&0&1.7188&14.0092&3.0413\\ 2.2528&5.6741&4.4252&3.5367&0&3.0413&30.9070\end{bmatrix},$
and the optimal decentralized feedback gain $K$ is given by
@@
The responses of all the state variables are illustrated in Figure 5, and the stability of system is guaranteed.
@@
$A=\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix},B_{1}=I_{3},B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
We utilize Algorithm 2 and Algorithm 3 respectively and the results are shown below (see Figure 6 and Figure 7).
@@
We can observe that the Algorithm 2 converges within 800 steps, while Algorithm 3 converges within only 80 steps.

### v3

@@ 6 Directly Optimizing $\ell_{0}$-penalty
$\min\limits_{\widetilde{W}}{}$ $u_{1,\sigma}(\widetilde{W},X_{k};\lambda)$ (62)
@@
with
$\mathcal{X}_{1}=\{\widetilde{W}\colon\widetilde{W}\in\Gamma_{+}^{p},\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},i=1,\dots,M,\widehat{A}\widetilde{W}=0\}$
which is closed convex set. First, we claim that $\{H_{\sigma}(\widetilde{W}_{k})\}$ is monotonically nonincreasing.
@@
Given an initial feasible point $\widetilde{W}_{0}$, let $\widetilde{W}_{k}$ be the iterative sequence generated by Algorithm 1 with $u_{1},u_{2}$ given by (61), (58). Then $H_{\sigma}(\widetilde{W}_{k+1})\leq H_{\sigma}(\widetilde{W}_{k}).$
Proof: Let $F(\widetilde{W},X_{k})=\langle{\rm vec}(R),\widetilde{W}\rangle+r_{2,\sigma}(y_{k})+h(\widetilde{W},y_{k})$. According to the fact that
@@
$H_{\sigma}(\widetilde{W}_{k+1})\leq u_{1,\sigma}(\widetilde{W}_{k+1},X_{k};\lambda)\leq u_{1,\sigma}(\widetilde{W}_{k},X_{k};\lambda)$
$=F(\widetilde{W}_{k},\widetilde{W}_{k})=H_{\sigma}(\widetilde{W}_{k}).$
[…]
Obviously, $\{H_{\sigma}(\widetilde{W}_{k})\}$ is bounded below by 00. Hence, by basic mathematic analysis, $H_{\sigma}(\widetilde{W}_{k})\downarrow H_{\sigma}^{\infty}$, where $H_{\sigma}^{\infty}$ is a positive constant. An inquiry that arises is whether or not $H_{\sigma}^{\infty}=H_{\sigma}^{*}\doteq\min H_{\sigma}(\widetilde{W})$. Regretfully, such question remains open, but we have the following result.
@@
Every cluster point $z=(z_{1},z_{2})$ of the iteration generated by Algorithm 1 with $u_{1},u_{2}$ given by (61), (58) is a coordinatewise minimizer of the optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$). In addition, if $H_{\sigma}(\cdot,\cdot)$ is regular at $z$, then $z$ is a stationary point of ($\textbf{P}^{\prime\prime}_{\sigma}$).
Proof: For a feasible $X_{0}=(\widetilde{W}_{0},y_{0})$, by (57), the sublevel set
@@
[…]
There exists a significant gap between coordinatewise minimizer and stationary point (or local minimizer). When the objective function is regular at coordinatewise minimizer, the coordinatewise minimum becomes a local minimum by the above theorem. In Lemma 3.1 of [59], there is a few discussion about regularity property of a Gateaux-differentiable function, but $H_{\sigma}(\widetilde{W},y)$ is generally not Gateaux-differentiable in this paper. Furthermore, if we assume that $H_{\sigma}^{\infty}=H_{\sigma}^{*}$ holds, we can obtain the following theorem.
@@
Proof: Since $f_{\sigma}(|(V_{2}\otimes V_{1})\widetilde{W}|)\geq 0$, it is obvious that, for all $\alpha\in\mathbb{R}$, ${\rm lev}_{\leq\alpha}H_{\sigma}(\widetilde{W})$ is compact. Then the set
@@
$\min_{\widetilde{W}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|(V_{2}\otimes V_{1})\widetilde{W}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (63)
${\rm s.t.}{}{}$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},~{}i=1,\dots,M,$
@@
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $\langle{\rm vec}(R),\widetilde{W}\rangle+\gamma y_{k}^{\top}\left|\widetilde{P}\right|+\frac{1}{2\lambda}\left\|\widetilde{W}-\widetilde{W}_{k}\right\|^{2}$ (64)
${\rm s.t.}{}{}$ $\widetilde{W}\in\Gamma^{p}_{+},$
$\widetilde{\Psi}_{i}\in\Gamma_{+}^{n},~{}i=1,\dots,M,$
@@
$\min_{\widetilde{W},\widetilde{P}}{}{}$ $f_{1}(\widetilde{W};\lambda)+f_{2}(\widetilde{W})+g(\widetilde{P})$ (65)
${\rm s.t.}{}{}$ $\mathcal{A}\widetilde{W}+\mathcal{B}\widetilde{P}=0.$
Interestingly, $g(\widetilde{P})$ is nothing but a weighted-$\ell_{1}$-norm with weight values given by $y_{k}$. Hence, compared to optimization problem (19), the only difference is the smooth and strongly convex term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ in $f_{1}(\widetilde{W};\lambda)$. Thus, we can utilize optimization scheme (21) by slightly modifying Algorithm 3 (just change $w_{ij}$ to $y_{k}^{(ij)}$, where $y_{k}^{(ij)}$ is the $(i,j)$-th element of $y_{k}$).
@@
  Given : Feasible initial point $X_{0}=(\widetilde{W}_{0},y_{0})\in\mathcal{X}_{1}\times\mathbb{R}_{+}^{mn},\lambda>0,\gamma>0,\sigma_{0}\leq 1,\alpha\in(0,1)$
  Result: $X_{k}=(\widetilde{W}_{k},y_{k})$
  for $i=1,2,3\dots$ do
     Set $\sigma_{i}=\alpha\sigma_{i-1}$
     Set $k=0$
     for $k=0,1,2,...$ do
        Let $k=k+1,i=(k\mod 2)+1$
        if $i==1$ then
           $\widetilde{W}_{k}\leftarrow$ALM based Decentralized LQ
           Set $y_{k}=y_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
        end if
        if $i==0$ then
           Update $y_{k}=\nabla f_{\sigma_{i}}\left(\left|(V_{2}\otimes V_{1})\widetilde{W}_{k}\right|\right)$
           Set $\widetilde{W}_{k}=\widetilde{W}_{k-1},X_{k}=(\widetilde{W}_{k},y_{k})$
        end if
        if Stopping Criterion==absent$=$ =True then
           Break
        end if
     end for
     Set $X_{0}=X_{k}$
  end for
@@
the iterative sequence of the algorithm BSUM can also converge to the coordinatewise minimizers of $H_{\sigma}(\widetilde{W},y)$. However, we introduce the proximal term $\frac{1}{2\lambda}\|\widetilde{W}-\widetilde{W}_{k}\|^{2}$ into $u_{1,\sigma}$, which appears to be meaningless or potentially resulting in a decelerated convergence of the BSUM algorithm. Empirically, the proximal term makes the objective of (65) be strongly convex and accelerate the solving for subproblem (65). Therefore, a trade-off exists between the convergence rate of the algorithm BSUM for optimization problem ($\textbf{P}^{\prime\prime}_{\sigma}$) and the convergence rate of (21) for subproblem (65).
@@
The method proposed in this section exhibits a significant level of complexity, and there exists no theoretical evidence to support the superiority of the method introduced in this section over Algorithm 3 (or Algorithm 4). Therefore, this section mostly focuses on theoretical discussions, particularly about the variational characteristics of the optimization problem induced by the $\ell_{0}$-penalty and the relation between $\ell_{1}$-penalty problem (15) and $\ell_{0}$-penalty problem. It is recommended that, for finding decentralized controllers, Algorithm 3 (or Algorithm 4) is employed.
@@ 7 Numerical Examples
$A=\begin{bmatrix}0.2220&0.9186&0.7659\\ 0.8707&0.4884&0.5184\\ 0.2067&0.6117&0.2968\end{bmatrix},B_{1}=I_{3},C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},$
$B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
By Algorithm 3, we can obtain the relation between $\gamma$ and $N$ (the number of 00 in feedback gain $K$; see Fig. 2).
$\gamma=0$ $\gamma=1$ $\gamma=5$ $\gamma=10$ $\gamma=20$ $\gamma=50$
$J(K)$ 1.92 3.51 5.78 6.56 7.20 7.91
$N$ 0 1 2 3 3 3
Table 1: Relation between $\gamma$ and $J(K)$
@@
Remarkably, if $(A,B_{2})$ is known, we have $\langle R,W\rangle=J(K)$ rather than an upper bound of $J(K)$; see Remark 1 above. The TABLE 1 illustrates the relation between $\gamma$ and LQ cost $J(K)$, and we can observe that $J(K)$ increases with the increase of $\gamma$. In particular, when $\gamma=0$, the problem studied in this example degenerates to classic centralized LQ problem, which can be solved by letting $\gamma\to 0$ and utilizing Algorithm 3 (or Algorithm 4; see Proposition 2) or solving Algebraic Riccati Equation. As illustrated in Fig. 2 and TABLE 1, the parameter $\gamma$ serves as the balance between LQ cost $J(K)$ and the level of sparsity of feedback gain $K$. Additionally, the LQ cost changes very little and the level of sparsity $N$ remains unchange when $\gamma$ is large enough; which implies if we only need a sufficiently sparse feedback gain $K$, we can simply let $\gamma$ sufficiently large.
Specifically, when $\gamma=10$, we can obtain the following result (see Fig. 3). Denoting
$W_{k}=\mathrm{vec}^{-1}(\widetilde{W}_{k})=\begin{bmatrix}W_{1,k}&W_{2,k}\\ W_{2,k}^{\top}&W_{3,k},\end{bmatrix},$
where $\{\widetilde{W}_{k}\}$ is the iteration sequence of Algorithm 3, we have that
$K_{k}=W_{2,k}^{\top}W_{1,k}^{-1}=\begin{bmatrix}K_{11,k}&K_{12,k}&K_{13,k}\\ K_{21,k}&K_{22,k}&K_{23,k}\end{bmatrix}$
converges to the stabilizing feedback gain by Theorem 1 and Corollary 2, and indeed the colored graphs in Fig. 3 indicate the convergence of $K_{ij,k}$ for $i=1,2$, $j=1,2,3$.
@@
The optimal decentralized feedback gain $K$ is given by
@@
The responses of all the state variables are illustrated in Fig. 4, and it can be seen that the stability of system is guaranteed.
@@
$A=\begin{bmatrix}0.3079&0.1879&0.1797&0.2935&0.6537\\ 0.5194&0.2695&0.5388&0.9624&0.5366\\ 0.7683&0.4962&0.2828&0.9132&0.9957\\ 0.7892&0.7391&0.7609&0.5682&0.1420\\ 0.8706&0.1950&0.2697&0.4855&0.9753\end{bmatrix},$
$B_{2}=\begin{bmatrix}0.6196&0.6414\\ 0.7205&0.9233\\ 0.2951&0.8887\\ 0.6001&0.6447\\ 0.7506&0.2956\end{bmatrix},$
@@
Letting $\gamma=10$, the optimal decentralized feedback gain $K$ is given by
@@
The responses of all the state variables are illustrated in Fig. 5, and the stability of system is guaranteed.
@@
$A=\begin{bmatrix}0&1&0\\ 0&0&1\\ 0&0&0\end{bmatrix},B_{1}=I_{3},B_{2}=\begin{bmatrix}0.9315&0.7939\\ 0.9722&0.1061\\ 0.5317&0.7750\end{bmatrix},$
$C=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{bmatrix},D=\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}.$
We utilize Algorithm 3 and Algorithm 4 with $\gamma=10$, respectively, and the results are shown below (see Fig. 6 and Fig. 7).
@@
By using Algorithms 3 and 4, denote the obtained feedback gain as $K_{1},K_{2}$, respectively; we have $J(K_{1})=8.81$ and $J(K_{2})=9.39$. Moreover, it holds $\|K_{1}\|_{0}=3$ and $\|K_{2}\|_{0}=4$. Generally, the obtained LQ cost $J(K_{1})$ is slightly smaller than the one by Algorithm 4; this is mainly caused by $g_{\mathcal{Q}}(x)\gg g(x)$ when $\|x\|$ is large. Additionally, the obtained feedback gains exhibit nearly the same level of sparsity by using Algorithms 3 and 4, however, the required iteration numbers could be really different. As shown in Fig. 6 and 7, we can observe that Algorithm 3 converges within 800 steps, while Algorithm 4 converges within only 80 steps. That is because the Algorithm 3 exhibits $\mathcal{O}(1/k)$ convergence rate, while Algorithm 4 exhibits $\mathcal{O}(1/k^{2})$ convergence rate.